B. Calculate the mass in grams of each of the following samples.

a.1.50x1018 atoms Ne

b.8.35x1021 mole3cules CH4


1. What is the percent composition of the following compounds?

a.Naphtalene(C10 H8)

b Sucrose (C12 H22 O11)

c Aluminom sulfate (Al2SO4)3

1
Expert's answer
2021-06-02T06:02:30-0400

B) a. 1.59 × 1018/ 6.023 × 1023

0.00000264 × 20.1797

= 0.0000533g

b). 8.35 × 1021/6.023 × 1023

= 0.0139g

1.a) Molar Mass of Naphthalene = 128.071

C = 12.0107 × 10 = 120.107

= 120.107/128.071× 100 = 93.78%

H = 8.0635/128.071 × 100 = 6.3%